# COSINE-100U: Upgrading the COSINE-100 Experiment for Enhanced Sensitivity to Low-Mass Dark Matter Detection

### v1

@@ VI Sensitivity of the COSINE-100U experiment
With the improved performance of higher light yields in the COSINE-100U detectors, along with background levels similar to those observed in the COSINE-100 experiment in the low-energy signal region, we evaluate the sensitivity of the COSINE-100U experiment for detecting dark matter, particularly for spin-dependent WIMP-proton interaction. We assume one year of operation, using the measured light yields at room temperature as summarized in table 1, and the expected background levels discussed in Section V, based on the COSINE-100 measurement Yu et al. (2024a). While the current COSINE-100 data analysis has reached an 8 NPE threshold Yu et al. (2024b), further improvements using machine learning techniques and simulated waveform of NaI(Tl) crystals Choi et al. (2024b) are expected to lower the threshold to 5 NPE, a level already achieved by the COHERENT experiment with CsI(Tl) crystal Akimov et al. (2017). For the sensitivity analysis of the COSINE-100U experiment, we assume a 5 NPE analysis threshold for each crystal.
@@
We generate WIMP interaction signals with and without the Midgal effect Migdal (1941); Ibe et al. (2018); Adhikari et al. (2022b). These signals are simulated for various interactions and masses within the standard WIMP galactic halo model Lewin and Smith (1996); Freese et al. (2013). Form factors and proton spin values of the nuclei are implemented using the publicly available dmdd package Gluscevic and McDermott (2015); Gluscevic et al. (2015); Anand et al. (2014); Fitzpatrick et al. (2013); Gresham and Zurek (2014). The electron-equivalent energy of the nuclear recoil is reduced using nuclear recoil quenching factors, which represent the ratio of scintillation light yield from sodium or iodine recoil relative to that from electron recoil for the same energy. Recently measured quenching factor values Lee et al. (2024b) were used, and the inclusion of the Migdal effect in the NaI(Tl) crystals follows our previous study Adhikari et al. (2022b). The measured electron-equivalent energy is then converted to NPE based on the light yield shown in Table 1.
@@
Poisson fluctuations in the measured NPE are considered for detector resolution, using a recently developed waveform simulation package Choi et al. (2024b), which has already provided a good description of energy resolution in the low-energy signal region of COSINE-100 data Lee et al. (2024a). We use an ensemble of simulated experiments to estimate the sensitivity of the COSINE-100U experiment, expressed as the expected cross-section limits for the WIMP-proton spin-dependent interactions in the absence of signals. For each experiment, a simulated spectrum is generated under a background-only hypothesis based on assumed background levels. Gaussian fluctuations of background components from the COSINE-100 measurement Yu et al. (2024a), along with COSINE-100U-specific background contributions discussed in Section V, and Poisson fluctuations in each energy bin are incorporated into each simulated experiment.
@@
We then fit the simulated data with a signal-plus-background hypothesis, applying flat priors for the signal and Gaussian constraints for the backgrounds. Systematic uncertainties affecting the background model are included as nuisance parameters Adhikari et al. (2021a). A Bayesian approach is used to analyze the single-hit energy spectrum between 5 NPE and 130 NPE for each WIMP model, covering several WIMP masses. Marginalization is performed to obtain the posterior probability density function for each simulated sample, allowing us to set the 90% confidence level exclusion limits. The 1,000 simulated experiments result in 68% and 95% bands of the expected limits, as shown in Fig. 17.
The expected limits from the COSINE-100U experiment are compared with the current best limits on low-mass WIMP-proton spin-dependent interactions from PICO-60 Amole et al. (2019), CRESST-III Li Angloher et al. (2022), NEWS-G Arora et al. (2024), and Collar Collar (2018). Leveraging the odd-proton numbers and relatively small atomic mass of sodium, the projected sensitivity of the COSINE-100U experiment can probe unexplored parameter spaces for WIMP masses below 4 GeV/c2, potentially reaching extremely low-mass regions as low as 20 MeV/c2 when considering the Migdal effect.
@@ VII Summary
The COSINE-100U experiment represents a major upgrade from the COSINE-100 experiment, aimed at improving sensitivity to low-mass dark matter detection. After 6.4 years of successful operation at the Yangyang Underground Laboratory, the experiment was relocated to the newly constructed Yemilab in Korea, which provides a deeper underground environment with enhanced shielding from cosmic muons. Key improvements in the COSINE-100U experiment include the implementation of a novel crystal encapsulation technique that increases light collection efficiency by approximately 35%. We have evaluated the expected sensitivities of the COSINE-100U experiment, assuming a total mass of 99.1 kg, a 1-year operation period, and a 5 NPE energy threshold. Under these conditions, the COSINE-100U detector has the potential to explore previously uncharted parameter spaces for spin-dependent WIMP-proton interactions.

### v2

@@ VI Sensitivity of the COSINE-100U experiment
With the improved performance of higher light yields in the COSINE-100U detectors, along with background levels similar to those observed in the COSINE-100 experiment in the low-energy signal region, we evaluate the sensitivity of the COSINE-100U experiment for detecting dark matter, particularly for spin-dependent WIMP-proton interaction. We assume one year of operation, using the measured light yields at room temperature as summarized in Table 2, and the expected background levels discussed in Section V, based on the COSINE-100 measurement COSINE-100:2024ola .
The current COSINE-100 data analysis achieved an 8 NPE threshold Yu:2024qgq ; however, further improvements are expected through the use of machine learning techniques and simulated waveforms of NaI(Tl) crystals Choi:2024ziz . These advancements are expected to lower the analysis threshold to 5 NPE, a level already achieved by the COHERENT experiment with a CsI(Na) crystal COHERENT:2017ipa . For the sensitivity analysis of the COSINE-100U experiment, we assume a 5 NPE analysis threshold for each crystal. Additionally, we evaluate sensitivities for different energy thresholds, including 8 NPE for the COSINE-100U setup and 8 NPE for the COSINE-100 setup, to illustrate the substantial improvements provided by the upgraded configuration.
@@
We generate WIMP interaction signals with and without the Midgal effect Migdal ; Ibe:2017yqa ; COSINE-100:2021poy . These signals are simulated for various interactions and masses within the standard WIMP galactic halo model Lewin:1995rx ; Freese:2012xd . Form factors and proton spin values of the nuclei are implemented using the publicly available dmdd package dmdd ; Gluscevic:2015sqa ; Anand:2013yka ; Fitzpatrick:2012ix ; Gresham:2014vja . The electron-equivalent energy of the nuclear recoil is reduced using nuclear recoil quenching factors, which represent the ratio of scintillation light yield from sodium or iodine recoil relative to that from electron recoil for the same energy. Recently measured quenching factor values Lee:2024unz are used, and the inclusion of the Migdal effect in the NaI(Tl) crystals follows our previous study COSINE-100:2021poy . Figure 15(a) shows the expected signal rates for two benchmark WIMP masses (2 and 5 GeV/c2) in spin-dependent WIMP-proton scattering scenarios without the Migdal effect. The measured electron-equivalent energy is converted to NPE using the light yield data in Table 2 and nonproportionality measurement in Ref. COSINE-100:2024log .
Figure 15: Expected signal rates and sensitivities with different energy thresholds. (a) The expected signal rates for two WIMP masses (Mχ) of 2 and 5 GeV/c2 are shown for three energy thresholds: COSINE-100 (8 NPE), COSINE-100U (8 NPE), and COSINE-100U (5 NPE), using the C6 crystal. (b) The expected 90% confidence level limits for the WIMP-proton spin-dependent cross-section are shown, assuming a 99.1 kg NaI(Tl) target operated for one year. The limits are presented for three different energy thresholds: COSINE-100 with 8 NPE, COSINE-100U with 8 NPE, and COSINE-100U with 5 NPE.
@@
Poisson fluctuations in the measured NPE are considered for detector resolution, using a waveform simulation package Choi:2024ziz , which has been validated with the low-energy signal region of COSINE-100 data COSINE-100:2024log . We use an ensemble of simulated experiments to estimate the sensitivity of the COSINE-100U experiment, expressed as the expected cross-section limits for the WIMP-proton spin-dependent interactions in the absence of signals. For each experiment, a simulated spectrum is generated under a background-only hypothesis based on assumed background levels. Gaussian fluctuations of background components from the COSINE-100 measurement COSINE-100:2024ola , along with COSINE-100U-specific background contributions discussed in Section V, and Poisson fluctuations in each energy bin are incorporated into each simulated experiment.
We then fit the simulated data with a signal-plus-background hypothesis, applying flat priors for the signal and Gaussian constraints for the backgrounds. Systematic uncertainties affecting the background model are included as nuisance parameters COSINE-100:2021xqn . A Bayesian approach is used to analyze the single-hit energy spectrum between 5 NPE (or 8 NPE) and 130 NPE for each WIMP model, covering several WIMP masses. Marginalization was performed to obtain the posterior probability density function for each simulated sample, allowing us to set the 90% confidence level exclusion limits.
To evaluate performance improvements, we consider three scenarios for detector performance, as shown in Fig. 15. The improved COSINE-100U encapsulation achieves approximately 80 times better sensitivity for a WIMP mass of 2 GeV/c2, assuming the same 8 NPE threshold as in the COSINE-100 analysis Yu:2024qgq . Lowering the threshold to 5 NPE further improves sensitivity by approximately 14 times for a 2 GeV/c2 WIMP mass.
The COSINE-100U expected limits, assuming a 5 NPE threshold and the measured light yields of Table 2 are compared with the current best limits on low-mass WIMP-proton spin-dependent interactions from PICO-60 PICO:2019vsc , CRESST-III Li CRESST:2022dtl , NEWS-G NEWS-G:2024jms , and Collar Collar:2018ydf as shown in Fig. 16. They are also compared to COSINE-100 limits from three years of data COSINE-100:2025xqn . Due to sodium’s odd-proton numbers and relatively low atomic mass, the COSINE-100U experiment has the potential to explore uncharted parameter spaces for WIMP masses below 3 GeV/c2, potentially reaching masses as low as 20 MeV/c2 when considering the Migdal effect.
@@ VII Summary
The COSINE-100U experiment represents a major upgrade from the COSINE-100 experiment, aimed at improving sensitivity to low-mass dark matter detection. After 6.4 years of successful operation at the Yangyang Underground Laboratory, the experiment was relocated to the newly constructed Yemilab in Korea, which provides a deeper underground environment with enhanced shielding from cosmic muons. Key improvements in the COSINE-100U experiment include the implementation of an improved crystal encapsulation technique that increases light collection efficiency by approximately 35%. We have evaluated the expected sensitivities of the COSINE-100U experiment, assuming a total mass of 99.1 kg, a 1-year operation period, and a 5 NPE analysis energy threshold. Under these conditions, the COSINE-100U detector has the potential to explore previously uncharted parameter spaces for spin-dependent WIMP-proton interactions.